The manufacturer of a certain kind of light bulb claims that the lifetime of the bulb in hours, X can be modeled as a random quantity with PDF.



fX(x)=0,c/x^2



x< 100



x> 100, where e is a constant. i. What value must e take in order for this to define a valid PDF? ii. What is the probability that the bulb lasts no longer than 150 hours? ii. Given that a bulb lasts longer than 150 hours, what is the probability that it lasts longer than 200 hours? Evaluate the cumulative distribution function.

1
Expert's answer
2022-02-11T12:51:57-0500

"i)"

"\int_{-\infty}^{\infty} f(x)dx=c\int_{100}^{\infty}{\frac 1 {x^2}}dx=-c\times{\frac 1 x}|_{100}^{\infty}={\frac c {100}}=1\implies c=100"



"ii)"

(ii) "P(X<150)=100\int_{100}^{150}{\frac 1 {x^2}}dx=-100\times{\frac 1 x}|_{100}^{150}=-100({\frac 1 {150}}-{\frac 1 {100}})={\frac 1 3}"


"iii)"

 "P(X<200|X>150)={\frac {P(150<X<200)} {P(X>150)}}"

"P(X>150)=1-P(X<150)=1-{1\over3}={\frac 2 3}"

"P(150<X<200)=100\int_{150}^{200}{\frac 1 {x^2}}dx=-100\times{\frac 1 x}|_{150}^{200}=-100\times({\frac 1 {200}}-{\frac 1 {150}})={\frac 1 6}"

So,

 "P(X<200|X>150)={\frac {{\frac 1 6}} {\frac 2 3}}={1\over4}"


"iv)"

Let F(x) be a cumulative distribution function, then "F(x)=\int_{-\infty}^x f(t)dt".

So, "F(x) = 0" when "x < 100" and,

"F(x)=100\u222b ^{\nx}_\n{100}\n\u200b{1\\over\n \nt^ \n2}\n \n\u200b\n dt=\u2212100\\times{1\\over \nt\n}\n\u200b\n \u2223 ^{\nx}_\n{100}\n\u200b\n =1\u2212 \n{100\\over x}\n\u200b" when "x \ge100"